Olive2006:

Could someone help with this please: 3) In the proportion a/b = c/d, what are the extremes and what are the means?

Vocaloid:

This is a definition-based question. I’m proportions of the format a/b = c/d, the extremes are the values at either end (the first and the last values, so a and d) The values in the middle (b and c) are the means.

Olive2006:

@vocaloid wrote:
This is a definition-based question. I’m proportions of the format a/b = c/d, the extremes are the values at either end (the first and the last values, so a and d) The values in the middle (b and c) are the means.
thank you this helps a lot